The diagonal of a computer monitor is measured to be 17 inches. If the width of the monitor is 12 inches, find the height of the monitor to the nearest inch. Show all work.
Answer by checkley77(12844)

A^2+B^2=C^2
12^2+B^2=17^2
144+B^2=289
B^2=289-144
B^2=145
B=SQRT145
B=12 INCHES.